rite the slope-intercept form of the equation of the line tha passes through (-2,4) and is perpendicular to the line 2x+ 3y=17
Answer by ewatrrr(24785)

 
Hi,


line passing through (-2,4) and perpendicular to the line:
2x+ 3y=17 OR y = -(2/3)x + 17/3 ||slope m = (-2/3)
New Line: m = 3/2 (negative reciprocal of (-2/3)
y = (3/2)x + b ||Using Pt(-2,4) to solve for b
4 = (3/2)· (-2) + b
7 = b
y = (3/2)x + 7